OpenStudy (anonymous):

Heather's family bought tickets at a local carnival. Her brother took 1/5 the number of tickets that her cousin took, which was half the number Heather took. If her brother took 10 tickets, how many tickets did Heather take?

OpenStudy (zpupster):

what we know: brother took 10 tickets---------b=10 brother took 1/5 of cousins-------1/5c=b or c=(5)(10) or c =50 that was 1/2 of Heather's----- 1/2h = c or h =(2)(50) or h=100

OpenStudy (anonymous):

yes